9. Determine the flow rate of IV administration of 1200ml to be given at a rate of 3mL/minute. The drop factor is 15.
Answer #1

3,l/mim= 180ml/hr

for 1200ml infusion time required will be= 1200/180= 15 hr= 900min

flow rate in gtt/min= (total volume/time in min) X drop factor in gtt/ml

=(1200/900) X15

= 19.99= 20 gtts/min